### v1

@@ 3. The Rank-3 Principal Stratum Automaton $\widehat{\mathcal{A}_{3}}$
Using the notation
@@
(6) $f_{k}^{(j)}:=g_{\sigma_{k-j-1}\circ\cdots\circ\sigma_{0}}^{-1}\circ f_{k}\circ g_{\sigma_{k-j-1}\circ\cdots\circ\sigma_{0}}\colon\Gamma_{k-1}^{(j)}\to\Gamma_{k}^{(j+1)},$
we have the following commutative diagram indicating how the map of (4) can be rewritten as folds and then possibly a single ltt structure permutation relabeling at the end.
@@
If necessary, we can increase $p$ so that $\sigma^{p}$ is the identity. Since $g^{p}$ can be decomposed as
@@
Since $g^{p}$ must also be principal, hence lone axis, this is the unique Stallings fold decomposition for $g^{p}$, so must be partial-fold conjugate to a sequence represented by a directed loop in $\mathcal{A}_{3}$.
@@
composes to be the same map as (8). We can see that $\Gamma_{1}$ has a valence-4 vertex as follows. Complete the fold
[Uncaptioned image]of the illegal turn at $\Gamma_{0}$. If it fully identifies two edges (upper right-hand image), the terminal vertex would have valence-5 and neither $g$ nor its fold-conjugates could have represented a principal $\varphi\in\textup{Out}(F_{3})$, a contradiction. So $g_{1}$ must be a proper full fold (lower right-hand image) and $\Gamma_{1}$ now has a valence-4 vertex at the image of $w$. So the previous paragraph indicates that
$\Gamma_{1}\xrightarrow{g_{2}}\Gamma_{2}\xrightarrow{g_{3}}\cdots\xrightarrow{g_{t-2}}\Gamma_{t-1}\xrightarrow{g_{\sigma}}\sigma\cdot\Gamma_{t-1}\xrightarrow{g_{\sigma}\circ g_{t-1}\circ g_{\sigma}^{-1}}\sigma\cdot\Gamma_{t}=\Gamma_{0}\xrightarrow{g_{1}}\Gamma_{1}$
is represented by a directed loop in $\widehat{\mathcal{A}_{3}}$. So in this case also the Stallings fold decomposition of $g$ is partial-fold conjugate to one determining a directed loop in $\widehat{\mathcal{A}_{3}}$. ∎
@@
Every principal fully irreducible $\varphi\in\textup{Out}(F_{3})$ contains in its Stallings fold decomposition the left-hand graph. Thus, every principal axis must pass through the simplex with underlying graph both the second and third graphs in the image.
@@
[Uncaptioned image]The image to the right schematically depicts $\widehat{\mathcal{A}_{3}}$, with the folds changing the labeling on the ltt structure included as dashed. The numbers on the arrows indicate how many distinct maps connect the isomorphism classes of ltt structures. It is an interesting observation that, as drawn to the right and in Figure 3, $\widehat{\mathcal{A}_{3}}$ has a clockwise orientation.
One can then observe that, once Node I is removed the maximal strongly connected component, which we call here $\mathcal{M}_{I}$, includes none of Node I, Node II, or Node III. Further, in the edge-labeling scheme of Figure 3, no permutation within $\mathcal{M}_{I}$ contains $c$ and no fold within $\mathcal{M}_{I}$ maps $c$ over another edge.
Hence, no directed loop within $\mathcal{M}_{I}$ can define an irreducible map. And we thus have that any irreducible map represented by a directed loop in the automata would need to contain Node I, which is the left-hand image in the statement of the proposition. The underlying graph of this ltt structure is the middle image. Inspection of all 4 folds entering Node I (including the self-maps), indicates that the axis must also pass through the third graph in the image. ∎
@@ 4. The single-fold map
Figure 2. The Stallings fold decomposition of $\mathfrak{g}$ is a fold $f$, then homeomorphism $g_{\sigma}$.
@@
Since the only illegal turn $\{d,\overline{c}\}$ is not taken, $\mathfrak{g}$ is a tt map. Since an adequately high power of the transition map is positive, $\mathfrak{g}$ is Perron–Frobenius. Alternatively, one can note that $\mathfrak{g}^{13}(d)$ contains all edges of $\Gamma$ and $d$ is in $\mathfrak{g}^{2}(a)$, $\mathfrak{g}(b)$, $\mathfrak{g}^{4}(c)$, and $\mathfrak{g}^{3}(e)$. So that all edges of $\Gamma$ are in $\mathfrak{g}^{15}(a)$, $\mathfrak{g}^{16}(b)$, $\mathfrak{g}^{17}(c)$, and $\mathfrak{g}^{16}(e)$. So that $\mathfrak{g}^{17}$ maps every edge over every edge.
In light of (12), the local Whitehead graphs are triangles at each valence-3 vertex, and then the union of the Figure 2 colored edges at the valence-4 vertex. Hence, each local Whitehead graph is connected. Using the sage package of Coulbois [Cou14], for example, one can check that there are indeed no PNPs.
@@ 5. Perron numbers & Minimal stretch factors
Suppose that $\varphi\in\textup{Out}(F_{r})$ is a fully irreducible outer automorphism. Its stretch factor is the PF eigenvalue $\lambda(g)$ of the transition matrix $M(g)$ for any expanding irreducible train track map $g$. By definition, $M(g)$ is a nonnegative integer matrix. By [Kap14, Lemma 2.4(2)], $M(g)$ is primitive. Hence, $M(g)$ is aperiodic. Since Perron numbers are precisely the spectral radii of nonnegative aperiodic integral matrices, $\lambda(g)$ is a Perron number, as desired. ∎
@@
$M(\mathfrak{g})=\begin{bmatrix}0&0&0&0&1\\ 1&0&0&0&0\\ 0&0&0&1&0\\ 0&1&0&0&0\\ 0&0&1&1&0\end{bmatrix}$
@@
Graphs in $CV_{3}$ have between 3 and 6 edges. Suppose $\varphi\in\textup{Out}(F_{3})$ has a tt representative on a graph with 6 edges. By an Euler characteristic argument, each graph with 6 edges is trivalent. If $\varphi$ is fully irreducible, then its Stallings fold decomposition must consist of at least one fold. As in the proof of Theorem C (and its corresponding image), the graph at the completion of the fold will either have a valence-4 vertex (if it is a proper full fold) or a valence-5 vertex (if it completely identifies 2 edges). In either case an Euler characteristic argument indicates the number of edges would have decreased from 6. Now, fold-conjugate outer automorphisms are in the same conjugacy class, so have the same stretch factor. Thus, if, a fully irreducible stretch factor is achieved by a tt map on a 6-edge graph, it is also achieved by a train track map on a graph with fewer edges. In fact, if a given stretch factor is achieved by a principal fully irreducible outer automorphism with a tt map on a 6-edge graph, then the same can be said on some graph of fewer edges.
In light of the previous paragraph, since the stretch factor of any element of $\textup{Out}(F_{3})$ is the largest eigenvalue of the transition matrix of a tt representative, the algebraic degree of the stretch factor is between 1 and 5. Moreover, the stretch factor must be a Perron number by Lemma 5.1. The smallest few Perron numbers of degrees 2, 3, 4, and 5 are known. Approximately, these are 1.618, 1.325, 1.221, and 1.124 respectively. ([Boy85, Table 3]) Since $\mathfrak{g}$ has stretch factor less than 1.32 and 1.22, the smallest stretch factor must be a degree-5 Perron number.
@@ 6. Uniqueness of the single-fold principal train track map
[Uncaptioned image]The goal of this section is to prove Theorem B. Assume throughout that $\Gamma$ is a graph of rank $r\geq 3$. Suppose that $h=g_{\sigma}\circ f$ is a tt representative of a principal fully irreducible $\varphi\in\textup{Out}(F_{r})$, such that $f\colon\Gamma\to\Gamma^{\prime}$ is a single proper full fold defined by $f\colon e_{1}\mapsto e_{0}^{\prime}e_{1}^{\prime}$ and $g_{\sigma}\colon\Gamma^{\prime}\to\Gamma$ is a graph relabeling isomorphism. Notice that one could not have $h=g_{\sigma}\circ f$ if $f$ were instead a partial fold, because partial folds change the graph-isomorphism type of a graph.
@@
Suppose $\Gamma$ is a graph with rank $r$ and $h=g_{\sigma}\circ f$ is an irreducible train track representative of a principal fully irreducible $\varphi\in\textup{Out}(F_{r})$ such that $f\colon\Gamma\to\Gamma^{\prime}$ is a single proper full fold defined by $f\colon e_{1}\mapsto e_{0}^{\prime}e_{1}^{\prime}$ and $g_{\sigma}\colon\Gamma^{\prime}\to\Gamma$ is a graph relabeling isomorphism, as described above. Then
@@
    $g_{\sigma}(v_{1}^{\prime})=v_{0}$, $g_{\sigma}(e_{1}^{\prime})=e_{0}$, and $g_{\sigma}(v_{2}^{\prime})=v_{1}$
@@
    Since the rank of $\Gamma$ is $r$, the Euler characteristic satisfies $\chi(\Gamma)=r-1$. So
    $r-1=\,|\,V\Gamma\,|\,-\,|\,E\Gamma\,|\,.$
    Since $\Gamma$ has a single valence-4 vertex and all other vertices are valence-3, we have:
@@
    In light of (b) we have the following picture (and will show the vertex colors reflect their orbit):
    Since $v_{0}$ is the unique valence-4 vertex of $\Gamma$, $v_{1}^{\prime}$ is the unique valence-4 vertex of $\Gamma^{\prime}$, and $g_{\sigma}$ is a graph isomorphism, we must have $g_{\sigma}(v_{1}^{\prime})=v_{0}$.
@@
    for all powers $k\geq 1$. Observe that $e_{1}^{\prime}$ is the single direction in $\Gamma^{\prime}$ not in the image of $Df$.
    We claim $Dg_{\sigma}(e_{1}^{\prime})\in\{e_{0},e_{1}\}$. Suppose not. Since $g_{\sigma}$ is a graph isomorphism, $Dg_{\sigma}$ is bijective on the set of directions of $\Gamma^{\prime}$. Hence there exists a pair of distinct directions $a^{\prime},b^{\prime}\in\mathcal{D}\Gamma^{\prime}$ such that $Dg_{\sigma}(a^{\prime})=e_{1}$ and $Dg_{\sigma}(b^{\prime})=e_{0}$. Since $g_{\sigma}(v_{1}^{\prime})=v_{0}$ and $g_{\sigma}$ is a graph-isomorphism, $a^{\prime}$ and $b^{\prime}$ emanate from $v_{1}^{\prime}$. Since we assumed neither $a^{\prime}$ nor $b^{\prime}$ are equal to the $Df$-unachieved direction $e_{1}^{\prime}$, the definition of $\Gamma^{\prime}$ ensures there exist $a,b\in\mathcal{D}\Gamma$ so that $Df(a)=a^{\prime}$ and $Df(b)=b^{\prime}$. Further, since $a^{\prime}$ and $b^{\prime}$ are incident to $v_{1}^{\prime}$ and $f(v_{1})=v_{1}^{\prime}$, we have that $a$ and $b$ are incident to $v_{1}$. Thus,
@@
    Since $g_{\sigma}$ is a graph isomorphism, the previous paragraph implies that $g_{\sigma}(e_{1}^{\prime})\in\{e_{1},e_{0}\}$. Suppose for the sake of contradiction that $g_{\sigma}(e_{1}^{\prime})=e_{1}$. Then $g_{\sigma}(v_{2}^{\prime})=v_{2}$, and so $h$ fixes the vertex $v_{2}$. Since $Dh(\overline{e_{1}})=Dg_{\sigma}(Df(e_{0}^{\prime}))=Dg_{\sigma}(\overline{e_{1}}^{\prime})=\overline{e_{1}}^{\prime}$, the set consisting of the remaining two edges incident to $v_{2}$ is invariant under $h$. This contradicts the irreducibility of $h$. Therefore $g_{\sigma}(e_{1}^{\prime})=e_{0}$.
@@
    Let $e\in E_{Y}$. By definition of $E_{Y}$, there is a vertex $v\in Y$ incident to $e$. Since $e_{1}\notin E_{Y}$, we know $e\neq e_{1}$, so by definition of $\Gamma^{\prime}$, we have $e^{\prime}$ is incident to $v^{\prime}$ in $\Gamma^{\prime}$. By definition of $f$, we have
@@
    Since $g_{\sigma}$ is a graph isomorphism, $g_{\sigma}(e^{\prime})$ remains incident to $g_{\sigma}(v^{\prime})$. Moreover, $g_{\sigma}(v^{\prime})=h(v)\in Y$ by invariance of $Y$. Hence $h(e)=g_{\sigma}(e^{\prime})\in E_{Y}$, so $E_{Y}$ is invariant under $h$, completing the proof of part (d). ∎
@@
Up to edge relabeling, the map $\mathfrak{g}$ of Figure 2 is the only train track map representing a principal fully irreducible outer automorphism whose Stallings fold decomposition consists of only a single fold composed with a graph-relabeling isomorphism.
@@
Suppose $h$ is a tt representative of a principal fully irreducible $\varphi\in\textup{Out}(F_{3})$. Then, by Proposition 3.4, $h$ is represented by a loop in $\widehat{\mathcal{A}_{3}}$ including the ltt structure (let us call it $G$) at Node I. If $h$ contains only a single fold, then this fold must be one of the folds taking $G$ to itself. Up to permutation-relabeling of ltt structures, there are only 2 such folds and the other combination of a fold then graph-relabeling isomorphism leaves invariant the subgraph labeled with $b$ and $c$ in the graph-labeling class of $\widehat{\mathcal{A}_{3}}$ included in Figure 3. Thus, up to edge relabeling, $h$ is unique.
@@
We maintain the convention that a vertex $v$ in $\Gamma$ is given the label $v^{\prime}$ in $\Gamma^{\prime}$, so we have a labeling of vertices in $\Gamma^{\prime}$, as well. Since $f(v)=v^{\prime}$ for any $v\in V\Gamma$, the above labeling implies $g_{\sigma}(v_{k}^{\prime})=v_{k-1}$ and $g_{\sigma}^{-1}(v_{k})=v^{\prime}_{k+1}$. Observe that by Lemma  6.1(c), our labeling is consistent with our original labeling of $v_{1}$ and $v_{2}$.
We will reach a contradiction to the existence of $h$ by showing $v_{2}$ would have the wrong valence. In what follows, we rely on the fact that the edge-labelings induced by $f$ (in the sense of §2.6) provides a bijection between $E\Gamma\backslash\{e_{1}\}$ and $E\Gamma^{\prime}\backslash\{[v_{1}^{\prime},v_{2}^{\prime}]\}$ given by mapping $e=[v,w]$ to $e^{\prime}=[v,w]$. Thus, if $e_{j}^{\prime}$ and $e_{k}^{\prime}$ have distinct vertices and are not $e_{1}^{\prime}=[v_{1}^{\prime},v_{2}^{\prime}]$ or $\overline{e_{1}^{\prime}}=[v_{2}^{\prime},v_{1}^{\prime}]$, then removing primes on vertices gives distinct edges of $\Gamma$.
@@
$\alpha_{1}=g_{\sigma}(e_{0}^{\prime})=[g_{\sigma}(v_{0}^{\prime}),g_{\sigma}(v_{1}^{\prime})]=[v_{n-1},v_{0}]$
Since $n\geq 5$, we know $v_{n-1}\neq v_{2}$, and hence also $\alpha_{1}\notin\{e_{1},\overline{e_{1}}\}$. By the definition of $\Gamma^{\prime}$ and its labeling, there is an edge $\alpha_{1}^{\prime}=[v_{n-1}^{\prime},v_{0}^{\prime}]\in E\Gamma^{\prime}$. Similarly, let
@@
Again, by the definition of $\Gamma^{\prime}$, there is an edge $\alpha_{2}^{\prime}=[v_{n-2}^{\prime},v_{n-1}^{\prime}]\in E\Gamma^{\prime}$. Recursively define

### v2

@@ 3.3. Loops in $\widehat{\mathcal{A}_{3}}$
Fixing a labeling on $\Gamma_{0}^{\prime}$, we induce labelings on all subsequent $\Gamma_{i}$ and $\Gamma_{i}^{\prime}$. Using the notation described in Definition 3.1 we can then define:
@@
(6) $f_{k}^{(j)}:=g_{\sigma_{k-j-1}\circ\cdots\circ\sigma_{0}}^{-1}\circ f_{k}\circ g_{\sigma_{k-j-1}\circ\cdots\circ\sigma_{0}}\colon\Gamma_{k-1}^{(j)}\to\Gamma_{k}^{(j+1)}.$
We have the following commutative diagram indicating how the map of (4) can be rewritten as folds and then possibly a single ltt structure permutation relabeling.
@@
If necessary, we can increase $p$ so $\sigma^{p}$ is the identity. Since $g^{p}$ can be decomposed as
@@
Since $g^{p}$ must also be principal, hence lone axis, this is the unique Stallings fold decomposition for $g^{p}$, so must be partial-fold conjugate to a fold decomposition represented by a directed loop in $\mathcal{A}_{3}$.
@@
composes to be the same map as (8). We can see that $\Gamma_{1}$ has a valence-4 vertex as follows. Complete the fold of the illegal turn at $\Gamma_{0}$. If it fully identifies two edges
[Uncaptioned image](upper right-hand image), the terminal vertex would have valence-5 and neither $g$ nor its fold-conjugates could have represented a principal $\varphi\in\textup{Out}(F_{3})$, a contradiction. So $g_{1}$ must be a proper full fold (lower right-hand image) and $\Gamma_{1}$ now has a valence-4 vertex at the image of $w$. So the previous paragraph indicates that $\widehat{\mathcal{A}_{3}}$ contains a directed loop representing
$\Gamma_{1}\xrightarrow{g_{2}}\Gamma_{2}\xrightarrow{g_{3}}\cdots\xrightarrow{g_{t-2}}\Gamma_{t-1}\xrightarrow{g_{\sigma}}\sigma\cdot\Gamma_{t-1}\xrightarrow{g_{\sigma}\circ g_{t-1}\circ g_{\sigma}^{-1}}\sigma\cdot\Gamma_{t}=\Gamma_{0}\xrightarrow{g_{1}}\Gamma_{1}.$
So in this case also the Stallings fold decomposition of $g$ is partial-fold conjugate to one determining a directed loop in $\widehat{\mathcal{A}_{3}}$. ∎
@@
Each principal fully irreducible $\varphi\in\textup{Out}(F_{3})$ contains in its Stallings fold decomposition the left-hand graph. Thus, each principal axis must pass through the simplex with underlying graph both the second and third graphs in the image.
@@
[Uncaptioned image]The image to the right schematically depicts $\widehat{\mathcal{A}_{3}}$, with folds changing the labeling on the ltt structure included as dashed. In this figure we group together the nodes differing only by an edge-label permutation into a single node, indexed by roman numerals; this grouping is also shown in Figure 3. Numbers on the arrows indicate how many distinct maps connect the isomorphism classes of ltt structures. It is interesting to note that, as drawn here and in Figure 3, $\widehat{\mathcal{A}_{3}}$ has a clockwise orientation.
One can then observe that, once Node I is removed, the maximal strongly connected component, which we call here $\mathcal{M}_{I}$, includes none of Node I, Node II, or Node III. Further, in the edge-labeling scheme of Figure 3, no permutation within $\mathcal{M}_{I}$ contains $c$ and no fold within $\mathcal{M}_{I}$ maps $c$ over another edge.
Hence, no directed loop within $\mathcal{M}_{I}$ can define an irreducible map. And we thus have that any irreducible map represented by a directed loop in the automaton would need to contain Node I, which is the left-hand image in the statement of the proposition. The underlying graph of this ltt structure is the middle image. Inspection of all 4 folds entering Node I (including the self-maps), indicates that the axis must also pass through the third graph in the image. ∎
@@ 4. The single-fold map
Figure 2. The Stallings fold decomposition of $\mathfrak{g}$ is a fold $f$, then homeomorphism $g_{\sigma}$. This map is indicated in dotted gold in 3.
@@
Since the only illegal turn $\{d,\overline{c}\}$ is not taken, $\mathfrak{g}$ is a tt map. Since an adequately high power of the transition map is positive, $\mathfrak{g}$ is Perron–Frobenius. Alternatively, one can note that $\mathfrak{g}^{13}(d)$ contains all edges of $\Gamma$ and $d$ is in $\mathfrak{g}^{2}(a)$, $\mathfrak{g}(b)$, $\mathfrak{g}^{4}(c)$, and $\mathfrak{g}^{3}(e)$. Thus all edges of $\Gamma$ are in $\mathfrak{g}^{15}(a)$, $\mathfrak{g}^{16}(b)$, $\mathfrak{g}^{17}(c)$, and $\mathfrak{g}^{16}(e)$. So $\mathfrak{g}^{17}$ maps every edge over every edge.
In light of (12), the local Whitehead graphs are triangles at each valence-3 vertex, and at the valence-4 vertex they are the union of the Figure 2 colored edges. Hence, each local Whitehead graph is connected. Using the sage package of Coulbois [Cou14], for example, one can check that there are indeed no PNPs.
@@ 5. Perron numbers & Minimal stretch factors
Suppose that $\varphi\in\textup{Out}(F_{r})$ is a fully irreducible outer automorphism. Its stretch factor is the PF eigenvalue $\lambda(g)$ of the transition matrix $M(g)$ for any expanding irreducible tt map $g$. By definition, $M(g)$ is a nonnegative integer matrix. By [Kap14, Lemma 2.4(2)], we know $M(g)$ is primitive. Hence, $M(g)$ is aperiodic. Since Perron numbers are precisely the spectral radii of nonnegative aperiodic integral matrices, $\lambda(g)$ is a Perron number, as desired. ∎
@@
$M(\mathfrak{g})=\begin{bmatrix}0&0&0&0&1\\ 1&0&0&0&0\\ 0&0&0&1&0\\ 0&1&0&0&0\\ 0&0&1&1&0\end{bmatrix}.$
@@
Graphs in $CV_{3}$ have between 3 and 6 edges. Suppose $\varphi\in\textup{Out}(F_{3})$ has a tt representative on a graph with 6 edges. By an Euler characteristic argument, each graph with 6 edges is trivalent. If $\varphi$ is fully irreducible, then its Stallings fold decomposition must consist of at least one fold. As in the proof of Theorem C (and its corresponding image), the graph at the completion of the fold will either have a valence-4 vertex (if it is a proper full fold) or a valence-5 vertex (if it completely identifies 2 edges). In either case an Euler characteristic argument indicates that the number of edges would have decreased from 6. Now, fold-conjugate outer automorphisms are in the same conjugacy class, so have the same stretch factor. Thus, if, a fully irreducible stretch factor is achieved by a tt map on a 6-edge graph, it is also achieved by a train track map on a graph with fewer edges. In fact, if a given stretch factor is achieved by a principal fully irreducible outer automorphism with a tt map on a 6-edge graph, then the same can be said on some graph of fewer edges.
In light of the previous paragraph, since the stretch factor of any element of $\textup{Out}(F_{3})$ is the largest eigenvalue of the transition matrix of a tt representative, the algebraic degree of the stretch factor is between 1 and 5. Moreover, the stretch factor must be a Perron number by Lemma 5.1. The smallest few Perron numbers of degrees 2, 3, 4, and 5 are known. Approximately, these are 1.618, 1.325, 1.221, and 1.124 respectively ([Boy85, Table 3]). Since $\mathfrak{g}$ has stretch factor less than 1.32 and 1.22, the smallest stretch factor must be a degree-5 Perron number.
@@ 6. Uniqueness of the single-fold principal train track map
[Uncaptioned image]The goal of this section is to prove Theorem B. Assume throughout that $\Gamma$ is a graph of rank $r\geq 3$. Suppose that $h=g_{\sigma}\circ f$ is a tt representative of a principal fully irreducible $\varphi\in\textup{Out}(F_{r})$ satisfying that $f\colon\Gamma\to\Gamma^{\prime}$ is a single proper full fold defined by $f\colon e_{1}\mapsto e_{0}^{\prime}e_{1}^{\prime}$ and $g_{\sigma}\colon\Gamma^{\prime}\to\Gamma$ is a graph relabeling isomorphism. Notice that one could not have $h=g_{\sigma}\circ f$ if $f$ were instead a partial fold or a complete fold, because these types of folds change the number of vertices in a graph, and hence change the graph-isomorphism type of a graph.
@@
Suppose $\Gamma$ is a graph with rank $r$ and $h=g_{\sigma}\circ f$ is an irreducible tt representative of a principal fully irreducible $\varphi\in\textup{Out}(F_{r})$ such that $f\colon\Gamma\to\Gamma^{\prime}$ is a single proper full fold defined by $f\colon e_{1}\mapsto e_{0}^{\prime}e_{1}^{\prime}$ and $g_{\sigma}\colon\Gamma^{\prime}\to\Gamma$ is a graph relabeling isomorphism. Label the vertices incident to the edges $e_{0},e_{1},e_{1}^{\prime}$, and $e_{0}^{\prime}$ as described above. Then
@@
    $g_{\sigma}(v_{1}^{\prime})=v_{0}$, $g_{\sigma}(e_{1}^{\prime})=e_{0}$, and $g_{\sigma}(v_{2}^{\prime})=v_{1}$, and
@@
    Since the rank of $\Gamma$ is $r$, the Euler characteristic satisfies $\chi(\Gamma)=1-r$. So
    $r-1=\,|\,E\Gamma\,|\,-\,|\,V\Gamma\,|\,.$
    Since $\Gamma$ has a single valence-4 vertex and all other vertices are of valence 3, we have:
@@
    In light of (b) we have the following picture (and will show the vertex colors reflect how they are mapped):
    Since $v_{0}$ is the unique valence-4 vertex of $\Gamma$ and $v_{1}^{\prime}$ is the unique valence-4 vertex of $\Gamma^{\prime}$ and $g_{\sigma}$ is a graph isomorphism, we must have $g_{\sigma}(v_{1}^{\prime})=v_{0}$.
@@
    for all powers $k\geq 1$. Observe that $e_{1}^{\prime}$ is the single direction in $\Gamma^{\prime}$ that is not in the image of $Df$.
    We claim $Dg_{\sigma}(e_{1}^{\prime})\in\{e_{0},e_{1}\}$. Suppose not. Since $g_{\sigma}$ is a graph isomorphism, $Dg_{\sigma}$ is bijective on the set of directions of $\Gamma^{\prime}$. Hence there exists a pair of distinct directions $a^{\prime},b^{\prime}\in\mathcal{D}\Gamma^{\prime}$ such that $Dg_{\sigma}(a^{\prime})=e_{1}$ and $Dg_{\sigma}(b^{\prime})=e_{0}$. Since $g_{\sigma}(v_{1}^{\prime})=v_{0}$ and $g_{\sigma}$ is a graph-isomorphism, $a^{\prime}$ and $b^{\prime}$ emanate from $v_{1}^{\prime}$. Since we assumed neither $a^{\prime}$ nor $b^{\prime}$ are equal to the $Df$-unachieved direction $e_{1}^{\prime}$, the definition of $\Gamma^{\prime}$ ensures there exists a pair of directions $a,b\in\mathcal{D}\Gamma$ so that $Df(a)=a^{\prime}$ and $Df(b)=b^{\prime}$. Further, since $a^{\prime}$ and $b^{\prime}$ are incident to $v_{1}^{\prime}$ and $f(v_{1})=v_{1}^{\prime}$, we have that $a$ and $b$ are incident to $v_{1}$. Thus,
@@
    Since $g_{\sigma}$ is a graph isomorphism, the previous paragraph implies that $g_{\sigma}(e_{1}^{\prime})\in\{e_{1},e_{0}\}$. Suppose for the sake of contradiction that $g_{\sigma}(e_{1}^{\prime})=e_{1}$. Then $g_{\sigma}(v_{2}^{\prime})=v_{2}$, and so $h$ fixes the vertex $v_{2}$. Since $Dh(\overline{e_{1}})=Dg_{\sigma}(Df(\overline{e_{1}}))=Dg_{\sigma}(\overline{e_{1}}^{\prime})=\overline{e_{1}}$, the set consisting of the remaining two edges incident to $v_{2}$ is invariant under $h$. This contradicts the irreducibility of $h$. Therefore $g_{\sigma}(e_{1}^{\prime})=e_{0}$.
@@
    Let $e\in E_{Y}$. By the definition of $E_{Y}$, there is a vertex $v\in Y$ incident to $e$. Since $e_{1}\notin E_{Y}$, we know $e\neq e_{1}$, so by the definition of $\Gamma^{\prime}$, we have $e^{\prime}$ is incident to $v^{\prime}$ in $\Gamma^{\prime}$. By the definition of $f$,
@@
    Since $g_{\sigma}$ is a graph isomorphism, $g_{\sigma}(e^{\prime})$ remains incident to $g_{\sigma}(v^{\prime})$. Moreover, $g_{\sigma}(v^{\prime})=h(v)\in Y$ by invariance of $Y$. Hence $h(e)=g_{\sigma}(e^{\prime})\in E_{Y}$, so $E_{Y}$ is $h$-invariant, completing the proof of (d). ∎
@@
Up to edge relabeling, the map $\mathfrak{g}$ of Figure 2 is the only train track map representing a principal fully irreducible outer automorphism in any rank whose Stallings fold decomposition consists of only a single fold composed with a graph-relabeling isomorphism.
@@
Suppose $h$ is a tt representative of a principal fully irreducible $\varphi\in\textup{Out}(F_{3})$. Then, by Proposition 3.4, $h$ is represented by a directed loop in $\widehat{\mathcal{A}_{3}}$ including the ltt structure (call it $G$) at Node I. If $h$ contains only a single fold, then this fold must be one of the folds taking $G$ to itself. Up to permutation-relabeling of ltt structures, there are only 2 such folds (the left-most 2 folds in the Node I box in Figure 3). $\mathfrak{g}$ is indicated using gold (dashed) arrows in Figure 3. In the graph-labeling class of $\widehat{\mathcal{A}_{3}}$ included in Figure 3, the other fold within the Node I box would lead to a fold of $d$ over $\bar{c}$ and then the graph isomorphism $a\mapsto\bar{e}$, $b\mapsto c$, $c\mapsto b$, $d\mapsto d$, $e\mapsto\bar{a}$ (a combination that leaves invariant $\{b\}\cup\{c\}$). Thus, up to edge relabeling, $h$ is unique.
@@
We maintain the convention that a vertex $v$ in $\Gamma$ is given the label $v^{\prime}$ in $\Gamma^{\prime}$, so we have a labeling of vertices in $\Gamma^{\prime}$ as well. Since $f(v)=v^{\prime}$ for each $v\in V\Gamma$, the above labeling implies $g_{\sigma}(v_{k}^{\prime})=v_{k-1}$ and $g_{\sigma}^{-1}(v_{k})=v^{\prime}_{k+1}$. Observe that by Lemma  6.1(c) our labeling is consistent with our original labeling of $v_{1}$ and $v_{2}$.
We will reach a contradiction to the existence of $h$ by showing $v_{2}$ would have the wrong valence. In what follows, we rely on the fact that the edge-labeling induced by $f$ (in the sense of §2.6) provides a bijection between $E\Gamma\backslash\{e_{1}\}$ and $E\Gamma^{\prime}\backslash\{[v_{1}^{\prime},v_{2}^{\prime}]\}$ given by mapping $e=[v,w]$ to $e^{\prime}=[v^{\prime},w^{\prime}]$. Thus, if $e_{j}^{\prime}$ and $e_{k}^{\prime}$ have distinct vertices and are not $e_{1}^{\prime}=[v_{1}^{\prime},v_{2}^{\prime}]$ or $\overline{e_{1}^{\prime}}=[v_{2}^{\prime},v_{1}^{\prime}]$, then removing primes on vertices gives distinct edges of $\Gamma$.
@@
$\alpha_{1}=g_{\sigma}(e_{0}^{\prime})=[g_{\sigma}(v_{0}^{\prime}),g_{\sigma}(v_{1}^{\prime})]=[v_{n-1},v_{0}].$
Since $n\geq 5$, we know $v_{n-1}\neq v_{2}$, and hence also $\alpha_{1}\notin\{e_{1},\overline{e_{1}}\}$. By the labeling of $\Gamma^{\prime}$, there is an edge $\alpha_{1}^{\prime}=[v_{n-1}^{\prime},v_{0}^{\prime}]\in E\Gamma^{\prime}$. Similarly, let
@@
Again, by $\Gamma^{\prime}$ labeling conventions, there is an edge $\alpha_{2}^{\prime}=[v_{n-2}^{\prime},v_{n-1}^{\prime}]\in E\Gamma^{\prime}$. Recursively define